Please show work that is not on excel please. Thank you. Рото 30. Systematic versus Unsystematic Risk. Consider the following information on Stocks I and II: State of Economy Probabi
Answer #1

a. Return of stock I = 22.4%
Return of stock II = 8.2%

b. Standard deviation of stock I = 12.73%
STANDARD DEVIATION OF STOCK II = 18.95%


c. 1. Capital Asset pricing model:

As per CAPM model:
Re= Rf+(Rm-Rf)B

Re= required rate of return. 22.4%
Rf= Risk-free rate. 4%
Rm =Market Risk Premium. 7%
B = Beta, systematic risk. ?

22.4 = 4+(7-4)B
Beta (B) = 6.13 (Stock I)

2. Capital Asset pricing model:

As per CAPM model:
Re= Rf+(Rm-Rf)B

Re= required rate of return. 8.2%
Rf= Risk-free rate (nominal) = 4%
Rm =Market Risk Premium. 7%
B = Beta, systematic risk. ?

8.2= 4+(7-4) x B
Beta = 1.4 (Stock II)


D. A has the most systematic risk as its beta is higher.
B has most unsystematic risk because it's differential between standard deviation and systematic risk is higher.
B is more risker because it has a higher standard deviation.

let Stock I> A Stock II=B Retuen P A B PA PB 0-25 24.- 204. 0.25x2 0.25x=20 O.GO 32% 12% 0.60×32 0.60x12 0.15 18%. 40% 0.15x1